# Harnessing quantum back-action for time-series processing

### v1

@@ IV.1 Performance analysis in the infinite resources scenario
$P_{R}(g,h)=\frac{C_{\Sigma}^{\text{OLP}}(g,h)}{\text{max}_{h}(C_{\Sigma}^{\text{RSP}}(h))}$ (6)
@@
We choose the best-performing RSP rather than the corresponding RSP performance for the same $h$, as we aim for a metric that indicates not only if the OLP performs as well as the RSP but also if the QRC algorithm performs well in general. While this is a strict metric, it avoids focusing on trivial scenarios where both protocols perform poorly. With this definition, a $P_{R}>1$ indicates a practical advantage of the OLP over the RSP. We average the $P_{R}$ over 50 different reservoir realizations, where the couplings in between the spins are randomly selected each time.
@@
In Figure 2 we show how $P_{R}$ changes at different $g$ and $h$ values for the forward prediction task on the Santa Fe dataset (top panel), and for the short-term memory task (bottom panel). Each colormap is made up of 12 by 40 points and is computed by assuming that we have access to an infinite amount of resources, i.e. an infinite number of shots for each measured observable. The $P_{R}$ at $g=0$ is set to be 00 since it corresponds to the case in which the measurement apparatus is not interacting with the reservoir and the measurement output is then a random number.
@@
Figure 2: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom). A $P_{R}$ in the range $[0,1)$ shows that the best performing RSP is performing better than the OLP in that point of the $g-h$ map. A $P_{R}$ around 1 shows that the OLP is performing as good as the best performing RSP, and over 1 indicates an over-performance.
@@
For both tasks, we can highlight the clear emergence of a region for which the OLP achieves comparable performances to the best-performing RSP. In the case of the forward prediction task, this region is spread along the $h$ axis while for the short-term memory task is peaked for $h$ close to $10^{-1}$. Moreover, by looking at the top panel we reveal a region in the parameter space where there is a numerical emergence of the OLP overperforming the RSP, i.e. a red region, for a $g$ value close to $1$. In this case, the back-action effect, which in principle is expected to negatively affect the performance of QRC, brings instead a positive contribution that allows the maximum performance of the RSP to be overcome.
@@
Hence, the emergence of these white and red regions highlights that a weak measurement-based protocol, if properly tuned, loses nothing in terms of performance compared to the RSP. To stress this last point, in Figure 3, we compare the different sum capacities of the QRC running with the RSP and with the OLP in the case in which $g$ is fixed to a non-optimal value and in the case in which it is fixed to the value that maximized $P_{R}$ in Figure 2. As a first observation, we can refer to [martinez2021dynamical] to justify the dependence of the performance on the $h$ value. In fact, along the horizontal axis, the quantum system undergoes a transition from a localized regime to an ergodic one and it was showcased that being in the ergodic regime is a preferential condition to perform reservoir computing. In addition, what emerges clearly from these plots is that by maximizing the $P_{R}$ one can optimize the strength of the measurement such that the reservoir performance is not negatively affected by the back-action and making the OLP as reliable as the RSP in terms of performance.
@@
The previous analysis is related to the sum capacity of the QRC across two different tasks. However, for different sub-tasks, defined by varying $\eta$, the performance may systematically differ. In this context, Figure 4 compares two points in the $g-h$ parameter space explored in Fig. 2, the one that maximized $P_{R}$ and a non-optimal point. We then evaluate the performance for each different sub-task relative to the performance of the RSP at optimal $h$.
Figure 4: Capacity for different sub-tasks. The figure compares the performance in single sub-tasks of the optimal restarting protocol with respect to the optimal and non-optimal versions of the OLP. The top panel shows the results for the forward prediction task while the bottom panel is for the short-term memory task.
@@ IV.2 Considering a finite amount of resources
In the previous section, we analyzed the effect of making use of a weak measurement-based approach to QRC in the ideal scenario of infinite resources. Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. In this context, it is important to consider the inherent inefficiency of the RSP compared to the OLP, as discussed in [mujal2023time]. This analysis takes into account the time resources required for executing each protocol. With the RSP, the time required to run the QRC scales quadratically with respect to the dimension of the time series, whereas for the OLP, this dependence is linear. Therefore, by fixing the number of shots available for the OLP, we can compute the required time and subsequently determine the number of shots available for the RSP within that same time frame. Approximately, we find that the number of shots available for the RSP is proportional to the number available for the OLP divided by $K$, the size of the dataset (see Appendix).
At this point, we can proceed to look at $P_{R}$, with the same logic as in Figure 2, but in a finite resources scenario. We now expect to see the combined effect of the back-action introduced by the OLP and of the higher efficiency of the OLP with respect to the RSP (Fig. 5). In this analysis, we fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$. Again the $P_{R}$ is probed in $12\times 40$ points in the $g-h$ parameter space, and each point is the average over 50 different reservoir realizations.
Figure 5: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom) in the scenario of finite resources. We fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$ and computed $N_{s}^{\text{RSP}}$ consequently for both tasks ($\approx 1.5\cdot 10^{3}$ for the forward prediction task and $\approx 3\cdot 10^{3}$ for the short-term memory task).
@@ V Conclusions and Discussion
In this work, we analyzed the effects of weak measurements on the QRC algorithm through a comprehensive comparison between the OLP and the RSP. Our findings reveal that, contrary to previous assumptions, the unwanted back-action from weak measurements can positively influence the performance of QRC algorithms. By optimizing the performance ratio $P_{R}$, we highlighted that an informed choice of measurement strength and external fields in the OLP allows overperform the commonly used RSP. The study is performed on several sub-tasks derived from two main benchmark tasks, widely used in the literature. Furthermore, we show that an optimal parameter choice for the OLP can be crucial to obtaining good performances depending on specific sub-tasks.
As the field transitions to experimental implementations, our results highlight the necessity of identifying optimal protocols, especially in the finite resources scenario. In this context, we combine the previous knowledge on the better scaling of the OLP compared to the RSP with the ability to positively affect the reservoir dynamics by tuning the measurement strength. Our findings highlight the clear advantage of the OLP in this setting. This practical orientation is essential as experimental interest in QRC continues to grow [Yelin2024, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, PhysRevApplied.20.014051], providing a clear roadmap for researchers aiming to harness the capabilities of quantum systems for data processing tasks.
@@
In conclusion, our analysis underscores the fundamental role of weak measurements in enhancing the performance of QRC through the OLP framework. By demonstrating that making use of weak measurements not only improves the scaling of the algorithm but can also improve its performance given an appropriate parameter selection, we lay the groundwork for future research.
Looking forward, we recognize the need for extended studies across a broader range of tasks, classical and quantum [Nokkala2023], and metrics, which may reveal even more generalized conclusions regarding the behavior of QRC algorithms under the OLP. In this direction, studying the effects of evolution time and coherence dynamics may provide deeper insights into the quantum processes underlying these systems. Our results strongly support the experimental realization of QRC algorithm with weak measurements through the OLP, particularly within emerging platforms that utilize neutral atoms and photons. As quantum computing evolves, these insights will drive both theoretical and practical applications in quantum machine learning.
@@ Acknowledgements
Details on the finite resource scenario
Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. Namely, for one and two spins observables the following statistical uncertainties are considered:
Within this scenario, it is also important to consider the inherent inefficiency of the RP compared to the OLP as discussed in [mujal2023time]. This analysis builds upon taking into account the time resources that each protocol requires to be executed. Let’s start by assuming that the time required to reset the reservoir and to take a measurement is negligible. Hence, we can define the experimental time required to run the QRC algorithm with the RSP on a dataset of $K$ elements as:

### v2

@@ IV.1 Performance analysis in the infinite resources scenario
$P_{R}(g,h)=\frac{C_{\Sigma}^{\text{OLP}}(g,h)}{\text{max}_{h^{\prime}}(C_{\Sigma}^{\text{RSP}}(h^{\prime}))}$ (7)
@@
We choose the best-performing RSP rather than the corresponding RSP performance for the same $h$, as we aim for a metric that indicates not only if the OLP performs as well as the RSP but also if the QRC algorithm performs well in general. While this is a strict metric, it avoids focusing on trivial scenarios where both protocols perform poorly. With this definition, a $P_{R}>1$ indicates a practical advantage of the OLP over the RSP. One could also choose to maximize $P_{R}$ separately for each direction, thereby obtaining optimal parameters independently along each axis. In Appendix B, we show that in our case this does not improve the results; however, it could guide an informed choice of the preferred measurement direction when resources are limited or bring advantages if one cares about a specific task with small $\eta$. We average the $P_{R}$ over 50 different reservoir realizations, where the couplings in between the spins are randomly selected each time. In the subsequent plots, the lines indicate the mean values, while the shaded regions denote one standard deviation around the mean.
@@
In Figure 2 we show how $P_{R}$ changes at different $g$ and $h$ values for the forward prediction task on the Santa Fe dataset (top panel), and for the short-term memory task (bottom panel). Each colormap is made up of 33 by 40 points and is computed by assuming that we have access to an infinite amount of resources, i.e. an infinite number of shots for each measured observable. The $P_{R}$ at $g=0$ is set to be 0 since it corresponds to the case in which the measurement apparatus is not interacting with the reservoir and the measurement output is then a random number uncorrelated to the system.
@@
Figure 2: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom). A $P_{R}$ in the range $[0,1)$ shows that the best performing RSP is performing better than the OLP in that point of the ghg-h map. A $P_{R}$ around 1 shows that the OLP is performing as good as the best performing RSP, and over 1 indicates an over-performance. The optimal parameters and corresponding $P_{R}$ values for each task are: forward prediction — RSP: $h^{\prime*}=0.346$; OLP: $g^{*}=0.355$, $h^{*}=0.273$, $P_{R}=1.006$; short-term memory — RSP: $h^{\prime*}=0.084$; OLP $g^{*}=0.256$ , $h^{*}=0.066$ , $P_{R}=1.029$.
@@
For both tasks, we can highlight the clear emergence of a region for which the OLP achieves comparable performances to the best-performing RSP. In the case of the forward prediction task, this region is spread along the $h$ axis while for the short-term memory task is peaked for $h$ close to $10^{-1}$. This indicates that the back-action effect, which would normally be expected to degrade the performance of QRC, instead provides a positive contribution that compensates for the limited information extracted from the system, enabling the OLP to match the performance of the RSP.
@@
Hence, the emergence of these white and almost-red regions highlights that an indirect measurement-based protocol, if properly tuned, loses nothing in terms of performance compared to the RSP. To stress this last point, in Figure 3, we compare the different sum capacities of the QRC running with the RSP and with the OLP in the case in which $g$ is fixed to a non-optimal value and in the case in which it is fixed to the value that maximized $P_{R}$ in Figure 2. As a first observation, we can refer to [martinez2021dynamical] to justify the dependence of the performance on the $h$ value. In fact, along the horizontal axis, the quantum system undergoes a transition from a localized regime to an ergodic one and it was showcased that being in the ergodic regime is a preferential condition to perform reservoir computing. Moreover, what emerges clearly is that optimizing the measurement strength, rather than simply minimizing back-action, is crucial for improving the protocol.
@@
The previous analysis is related to the sum capacity of the QRC across two different tasks. However, for different sub-tasks, defined byăvarying $\eta$, the performance may systematically differ. In this context, Figure 4 compares two points in the ghg-h parameter space explored in Fig. 2, the one that maximized $P_{R}$ and a non-optimal point. We then evaluate the performance for each different sub-task relative to the performance of the RSP at optimal $h$.
Figure 4: Capacity for different sub-tasks. The figure compares the performance in single sub-tasks of the optimal restarting protocol with respect to the optimal and non-optimal versions of the OLP. The top panel shows the results for the forward prediction task while the bottom panel is for the short-term memory task. The optimal parameters correspond to the values obtained in Figure 2, whereas the non-optimal parameters are fixed to $g=2$ and $h=38.88$.
@@ IV.2 Considering a finite amount of resources
In the previous section, we analyzed the effect of making use of an indirect measurement-based approach to QRC in the ideal scenario of infinite resources. Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time]. Namely, for one and two spins observables the following statistical uncertainties are considered:
For the RSP, one simply considers the asymptotic regime of very large $g$. In this context, it is important to consider the inherent inefficiency of the RSP compared to the OLP, as discussed in [mujal2023time]. This analysis takes into account the time resources required for executing each protocol. With the RSP, the time required to run the QRC scales quadratically with respect to the dimension of the time series, whereas for the OLP, this dependence is linear. Therefore, by fixing the number of shots available for the OLP, we can compute the required time and subsequently determine the number of shots available for the RSP within that same time frame. Approximately, we find that the number of shots available for the RSP is proportional to the number available for the OLP divided by $K$, the size of the dataset (see Appendix A).
At this point, we can proceed to look at $P_{R}$, with the same logic as in Figure 2, but in a finite resources scenario. We now expect to see the combined effect of the back-action introduced by the OLP and of the higher efficiency of the OLP with respect to the RSP (Fig. 5). In this analysis, we fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$. Again the $P_{R}$ is probed in $33\times 40$ points in the ghg-h parameter space, and each point is the average over 50 different reservoir realizations.
Figure 5: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom) in the scenario of finite resources. We fixed $N_{s}^{\text{OLP}}=1.5\cdot 10^{6}$ and computed $N_{s}^{\text{RSP}}$ consequently for both tasks ($\approx 1.5\cdot 10^{3}$ for the forward prediction task and $\approx 3\cdot 10^{3}$ for the short-term memory task). The optimal parameters and corresponding $P_{R}$ values for each task are: forward prediction — RSP: $h^{\prime*}=15.118$; OLP: $g^{*}=0.486$, $h^{*}=30.703$, $P_{R}=1.158$; short-term memory — RSP: $h^{\prime*}=1.245$; OLP $g^{*}=0.355$ , $h^{*}=0.170$ , $P_{R}=1.416$.
@@
IV.3 Comparison with a classical feedback-based protocol
So far we have compared the OLP driven by an optimal back-action with the restarting protocol. This allowed us to highlight the advantage of exploiting the information disturbance tradeoff with respect to an ideal but inefficient scenario were measurements do not affect the system. More recently, alternative online schemes have been proposed where the reservoir is restarted at every step but complemented with a classical feedback layer that reconstructs the reservoir state informed by the previous expectation values [Kobayashi2024]. This approach is undoubtedly practical and efficient, yet it relies entirely on a classical memory, which limits the exploitation of genuinely quantum resources. By contrast, in our indirect-measurement OLP protocol, although less information is extracted per measurement, the reservoir retains quantum coherences that constitute a form of quantum memory. We therefore expect that, at comparable resource scaling, the optimized OLP can outperform protocols that rely only on classical memory like the feedback-driven protocol proposed in [Kobayashi2024].
To test this expectation, we adapt the feedback-driven protocol to our spin-reservoir setting and benchmark it directly against our optimized OLP. The adapted protocol operates as follows. For each element $k$ of the input time series: (i) the value $s_{k}$ is encoded into one of the reservoir qubits with the same encoding map of the OLP; (ii) a unitary feedback layer is applied, which re-injects information about the previous expectation values $\langle\sigma^{i}_{k-1}\rangle$ into the reservoir; (iii) the system is evolved under the reservoir Hamiltonian (3); and (iv) all qubits are measured projectively and the reservoir is reset to its initial state before the next input. This routine is repeated sequentially for the entire time-series. To ensure consistency with our previous analysis we measure single-spin observables in all three directions ($x,y,z$), but exclude two-spin observables since the classical-feedback protocol was originally devised for single-spin readout. The feedback layer is implemented as a staggered array of two-qubit modules acting across the reservoir qubits (see Fig. 6). Each module is parameterized by the measured expectation values at the previous step and a feedback strength $a_{fb}$. This module acting on qubits $i$ and $j$ takes the form
$R_{ij}(\theta)=RX_{i}(\theta)RX_{j}(\theta)CX_{ij}RZ_{j}(\theta)CX_{ij},$ (9)
where $\mathrm{RX}(\cdot)$ and $\mathrm{RZ}(\cdot)$ denote single-qubit rotations, $\mathrm{CX}_{ij}$ is a CNOT with control $i$ and target $j$ and $\theta=a_{fb}\langle\sigma_{k-1}^{i}\rangle$.
Figure 6: Feedback-driven protocol. We adapt the feedback-driven protocol by inserting the unitary feedback layer in our reservoir pipeline. The measurements after the evolution are projective, and at each new input, we reset the system to the initial conditions.
We test this adapted feedback-driven model on the same short-term memory task considered above. As a first step, we vary the Hamiltonian parameter $h$, which controls the external field, and evaluate the sum capacity at different feedback strengths $a_{fb}$. The results (Fig. 7 (a)) show that the maximum sum capacity is obtained for $h=10$ and $a_{fb}$ $\approx 0.63$. We then fix $h=10$ and examine in detail the capacities of individual memory tasks at different delays $\eta$, while still scanning over $a_{fb}$. These are compared with the corresponding capacities of the optimally performing OLP (Fig. 7 (b)). The comparison reveals that at short delays ($\eta=1$) the optimal feedback-driven protocol performs comparably to the optimized OLP. However, as the task becomes more demanding (larger $\eta$), the advantage of the OLP becomes evident: the presence of quantum memory allows it to maintain higher performance, clearly surpassing the classical-memory-based feedback strategy.
Figure 7: Performance of the feedback-driven protocol across different systems (top) and versus the optimal OLP (bottom) for the short-term memory task. We consider an infinite resource scenario and each line is the average over 10 different realizations. The shaded regions denote one standard deviation around the mean. In the bottom panel, the system chosen for the feedback protocol has an $h$ value of 10.
@@ V Conclusions and Discussion
In this work, we analyzed the effects of indirect measurements on the QRC algorithm. Our findings reveal that, contrary to previous assumptions, the unwanted back-action from indirect measurements can positively influence the performance of QRC algorithms. By optimizing the performance ratio $P_{R}$, we highlighted that an informed choice of measurement strength and external fields in the OLP allows overperforming the standard RSP. The study is performed on several sub-tasks derived from two main benchmark tasks, widely used in the literature. We also show that an optimal parameter choice for the OLP can be crucial to obtaining good performances depending on specific sub-tasks. Moreover, we benchmark the optimal OLP against the feedback-driven protocol, a linear-scaling alternative that relies on classical memory and a unitary feedback layer. Our results demonstrate that the ability of the OLP to partially preserve coherences in the reservoir state provides an advantage that incoherent feedback alone cannot match.
As the field transitions to experimental implementations, our results highlight the necessity of identifying optimal protocols, especially in the finite resources scenario. In this context, we combine the previous knowledge on the better scaling of the OLP compared to the RSP with the ability to positively affect the reservoir dynamics by tuning the measurement strength. Together, these findings emphasize the OLP as a compelling candidate for near-term experimental demonstrations of QRC.This practical orientation is essential as experimental interest in QRC continues to grow [Yelin2024, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, PhysRevApplied.20.014051], providing a clear roadmap for researchers aiming to harness the capabilities of quantum systems for data processing tasks.
@@
In conclusion, our analysis underscores the fundamental role of indirect measurements in enhancing the performance of QRC through the OLP framework. By demonstrating that making use of indirect measurements not only improves the scaling of the algorithm but can also improve its performance given an appropriate parameter selection, we lay the groundwork for future research.
Looking forward, we recognize the need for extended studies across a broader range of tasks, classical and quantum [Nokkala2023], and metrics, which may reveal even more generalized conclusions regarding the behavior of QRC algorithms under the OLP. In this direction, studying the effects of evolution time and coherence dynamics may provide deeper insights into the quantum processes underlying these systems. Our results strongly support the experimental realization of QRC algorithm with indirect measurements through the OLP, particularly within emerging platforms that utilize neutral atoms and photons. As quantum computing evolves, these insights will drive both theoretical and practical applications in quantum machine learning.
@@ Acknowledgements
V.1 Details on the finite resource scenario
Here we consider the case in which each expectation value extracted from the reservoir is computed with a finite number of shots $N_{s}$. To take this into account, it has been shown that it is enough to add a Gaussian noise dependent on $N_{s}$ and $g$ to the previously computed expectation values [mujal2023time] (for further details, refer to Equation (8)).
Within this scenario, it is also important to consider the inherent inefficiency of the RSP compared to the OLP as discussed in [mujal2023time]. This analysis builds upon taking into account the time resources that each protocol requires to be executed. Let’s start by assuming that the time required to reset the reservoir and to take a measurement is negligible. Hence, we can define the experimental time required to run the QRC algorithm with the RSP on a dataset of $K$ elements as:
@@ V.1 Details on the finite resource scenario
V.2 Independent parameter optimization per direction
In this appendix we explore the effect of maximizing the performance ratio $P_{R}$ independently for each observable direction. We do this analysis for the short-memory task. Concretely, we consider three separate scenarios in which only the observables along a single axis (either $x$, $y$, or $z$) are used to compute $P_{R}$, and we determine the optimal parameters $(g,h)$ that maximize $P_{R}$ for that direction following the same procedure described in the main text. This per-direction analysis give intuition about the contribution of each direction to the total performance.
As shown in Figure 8, the $x$ direction contributes less to $P_{R}$ than the other two directions, which are very similar both in behaviour and in optimal parameters.
Figure 8: Performance ratio analysis for the short-term memory task, computed independently for each measurement direction. Each panel shows the ghg-h map when only observables in a single direction are used to evaluate $P_{R}$. The optimal parameters and corresponding $P_{R}$ values for each direction are: ($x$). RSP: $h^{\prime*}=0.170$; OLP : $g^{*}=0.150$, $h^{*}=0.170$, $P_{R}=0.981$ ($y$). RSP: $h^{\prime*}=0.084$; OLP: $g^{*}=0.355$, $h^{*}=0.041$, $P_{R}=1.145$ ($z$). RSP: $h^{\prime*}=0.084$; OLP: $g^{*}=0.355$, $h^{*}=0.041$, $P_{R}=1.159$.
Next, we take the observables corresponding to each direction’s independently optimized parameters and build a combined dataset used to solve the task. In Figure 9 we compare the capacities of this combined dataset with those obtained from the joint (global) optimization reported in the main text. The combined, per-direction strategy yields a small increase in the performance ratio, $P_{R}=1.040$ (vs. $P_{R}=1.029$ for the global optimisation). Although this increase has little impact on the summed capacity reported in the main text, the independent optimisation redistributes capacity from larger delays (the "tail") to shorter delays (the "head"): capacities at small delay $\eta$ are improved while long-$\eta$ performance is slightly reduced. Thus, for applications that prioritise short-delay memory, the per-direction strategy can be advantageous. Finally, since the $x$ direction contributes relatively little, one may reasonably omit $x$ observables when resources are constrained and focus on the more informative $y$ and $z$ directions.
Figure 9: Capacity for different sub-tasks. The figure compares the capacity of the OLP when optimized independently for each measurement direction and combined across directions with the OLP optimized jointly over all observables. The per-direction strategy gives a modest gain, redistributing capacity from long to short delays (see main text for discussion).